A 35 N force was applied to a 7-kg toy cart, causing it to accelerate to the right. Compute its acceleration.



1
Expert's answer
2021-10-22T02:05:26-0400

Given

F = 35 [N]

m = 7 [kg]

a - ?


Solution:

according to Newton's Second Law

a = F/m = 35/7 = 5 m/s^2